OpenStudy (anonymous):

Express the polynomial as a product of linear factors. ƒ(x)=3x^3+12x^2+3x-18 A. (x+3)(x+6)(x-1) B. 3(x-1)(x+3)(x+2) C. (x-2)(x+3)(x-3) D. (x-3)(x+3)(x-2)

myininaya (myininaya):

f(1)=0 so x=1 is a zero so x-1 is a factor

myininaya (myininaya):

1| 3 12 3 -18 | 3 15 18 --------------- 3 15 18| 0 so another factor is 3x^2+15x+18 but we can factor this factor

OpenStudy (anonymous):

Answer=B

myininaya (myininaya):

right you can also factor out 3 so we would have 3(x-1)(x^2+5x+6) =3(x-1)(x+2)(x+3)

OpenStudy (anonymous):

f(x)=3x^3+12x2+3x-18 the common factor is 3 - pull that out, you get 3(x^3+4x^2+x-6) so, 3(x-1)(x+2)(x+3) prove it... 3*2*-1=-6 3+2-1=4

OpenStudy (anonymous):

thanks!